# A Nonlocal Aw-Rascle-Zhang system with linear pressure term

### v1

@@ 3. The scalar balance law
for some given initial datum $M^{0}$. We can prove the following result.
@@
Assume (H) and (1.6). Consider the Cauchy problem (3.28) with $M^{0}$ a nondecreasing function. Suppose that there exists an $R^{0}>0$ such that
@@
    Given any $T>0$, the Cauchy problem (3.28) has a unique entropy solution
@@
    For any $T>0$, the entropy solution $M$ can be approximated by the discretized balance law (3.11) with initial datum (3.12), and thus by the sticky particle Cucker-Smale dynamics (2.1), i.e. there exists a sequence of initial data $M_{N}^{0}$ and fluxes $A_{N}$ satisfying the assumptions (3.14)-(3.13) respectively, such that the associated entropy solution $M_{N}$ constructed in Theorem 3.3 satisfies
@@
    Let $\tilde{M}$ be the entropy solution of
@@
[…]
@@
[…]
@@
Let $N\in\mathbb{N}$ be fixed, the aim is to construct two functions $M_{N}$ and $A_{N}$ such that they satisfy the assumptions of Theorem 3.3. Let us consider positive masses $m_{i,N}$ such that
@@
We can notice that $x_{i,N}^{0}\in[-R^{0},R^{0}],\,\forall i\in\{1,...,N\}$. At this point, we can construct the functions $M_{N}^{0}$ and $A_{N}$ as in the following:
@@
These functions obviously satisfy the hypotheses (3.14) and (3.13) of Theorem 3.3, thus we can write the entropy solution
@@
of the scalar balance law (3.11). The trajectories of the particles involved in the Cucker-Smale dynamics with the initial data $(m_{i,N},x_{i,N}^{0},v_{i,N}^{0})_{i=1}^{N}$ are $x_{i,N}(t)$ with $i\in\{1,...,N\}.$ The initial velocities can be obtained using the relation
@@
Moreover, the quantities $M_{N}^{0}$ and $A_{N}$ approximate the quantities $M^{0}$ and $A,$ i.e. the following inequalities hold true
[…]
@@
and it is possible to conclude that $M_{N}^{0}-M^{0}\to 0$ in $L^{1}(\mathbb{R})$ and $A_{N}\to A$ uniformly as $N\to\infty$ using (3.33).
@@
  • To prove [(3.35)], let us denote $x_{0,N}^{0}\mathrel{\mathop{\ordinarycolon}}=-R^{0}$ and we write
    $||M_{N}^{0}$ $-M^{0}||_{L^{1}(\mathbb{R})}=\sum_{j=1}^{N}\int_{x_{j-1,N}}^{x_{j,N}}|M_{N}^{0}(x)-M^{0}(x)|\,dx$
@@
Let $T>0$ be a fixed time, since $M_{N}(\cdot,t)$ is uniformly bounded and nondecreasing for any $t\in[0,T],$ we can apply the Helly’s Theorem stating that there exists a $L^{1}_{loc}(\mathbb{R})$-convergent subsequence $M_{N_{k}}(\cdot,t)$. Applying a diagonal argument, it is possible to get another subsequence convergent in $L^{1}_{loc}(\mathbb{R})$ for all rational $t\in[0,T]$, we denote again this subsequence as $M_{N_{k}}$ to simplify the notation and we denote with $M(\cdot,t)$ the limit of this subsequence. The aim is to extend this convergence from $L^{1}_{loc}(\mathbb{R})$ to $L^{1}(\mathbb{R})$ and also to extend the result to irrational times. From the sticky particle Cucker-Smale dynamics (2.5), we know $\{x_{i,N}(t)\}_{i=1}^{N}\subset[-R(T),R(T)]$ with $R(T)=R^{0}+T\tilde{M}$. This means that
@@
thus $M_{N_{k}}(t)-M(t)\to 0$ in $L^{1}(\mathbb{R})$ for all rational times $t\in[0,T]$.
@@
Moreover, by the time regular estimate (3.37), it holds
$\partial_{t}M_{N_{k}}(x,t)||_{\mathcal{M}}\leq\tilde{M},\qquad\forall t\in[0,T],$
@@
and for this reason we can conclude that $M\in BV(\mathbb{R}\times[0,T])$. Now we need to verify if the limit function $M$ satisfies the entropy condition (3.7). We check this using the Kružkov entropy pairs $(\eta,q)$ defined as in (3.6). We know that
@@
is an entropy solution of the discretized scalar balance law (3.11), this means that it satisfies the entropy inequality with the couple entropy/entropy-flux $(\eta,q_{N})$, where $\eta(m)=|m-\alpha|$ and $q_{N}(m)=\text{sgn}(m-\alpha)(A_{N}(m)-A_{N}(\alpha))$ with $\alpha\in[0,1]$, namely
@@
Now we consider the subsequence $M_{N_{k}}$ and we pass to the limit, to simplify the notation we use $M_{N}$ in place of $M_{N_{k}}$. For a fixed $t\in[0,T]$ we obtain
[…]
@@
About the last term, we start by showing that $\phi\ast M_{N}$ converges to $\phi\ast M$ as $N\to+\infty.$ Indeed, since $\phi\in L^{\infty}(\mathbb{R})$, then
[…]
@@
$\partial_{x}\eta(M_{N})||_{\mathcal{M}(\mathbb{R})}\leq|\eta|_{Lip}$
and this is also true for $\partial_{x}(\eta(M))$. Now, we are able to prove the convergence of the last term in the entropy inequality (3.39); indeed
@@
$\leq$ $\int_{0}^{T}||\varphi(t)||_{L^{\infty}(\mathbb{R})}||\phi\ast M_{N}-\phi\ast M||_{L^{\infty}(\mathbb{R})}||\partial_{x}(\eta(M_{N})||_{\mathcal{M}}\,dt$
$+\left|\int_{0}^{T}\int_{\mathbb{R}}(\phi\ast M)\varphi[\partial_{x}(\eta(M_{N}))-\partial_{x}(\eta(M))]\,dx\,dt\right|.$
Since $\phi\ast M_{N}-\phi\ast M||_{L^{\infty}(\mathbb{R})}\to 0$, then the first term on the right-hand side above converges to zero, while for the second term we have
$\left|\int_{0}^{T}\int_{\mathbb{R}}(\phi\ast M)\right$. $\left.-\varphi[\partial_{x}(\eta(M_{N}))-\partial_{x}(\eta(M))]\,dx\,dt\right|$
$=\left|\int_{0}^{T}\int_{\mathbb{R}}-\partial_{x}[(\phi\ast M)\varphi][(\eta(M_{N}))-(\eta(M))]\,dx\,dt\right|$
$\leq\int_{0}^{T}\int_{\mathbb{R}}\mathinner{\!\left\lvert-\partial_{x}[(\phi\ast M)\varphi][(\eta(M_{N}))-(\eta(M))]\right\rvert}\,dx\,dt$
$\leq\mathinner{\!\left\lVert\partial_{x}[(\phi\ast M)\varphi]\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{0}^{T}\mathinner{\!\left\lVert\eta(M_{N})-\eta(M)\right\rVert}_{L^{1}(\mathbb{R})}\,dt\,.$
Since $\eta$ is Lipschitz continuous and by means of (3.38), the right-hand side above converges to zero and thus, $M$ is an entropy solution of (3.28).
@@
$+\iint\text{sgn}(M(x,t)-\tilde{M}(y,s))(A(M(x,t)-A(\tilde{M}(y,s)))\partial_{x}w(x,t,y,s)\,dx\,dt$
$+\iint(\phi\ast M)(x,t)(\partial_{x}|M(x,t)-\tilde{M}(y,s)|\,dx\,dt.$
We perform the same for the entropy inequality of $\tilde{M}$, but with $\tilde{A}$ in place of $A$ and with $\alpha=\tilde{M}(y,s)$. We integrate both the inequalities above over the remaining free variables in both cases and add them, then we get
@@
$\iint\left[|M(x,t)-\tilde{M}(x,t)g(x)h^{\prime}_{\delta}(t)+\text{sgn}(M-\tilde{M})(A_{+}(M)-A_{+}(\tilde{M}))g^{\prime}(x)h_{\delta}(t)\right]\,dxdt$
@@
$+\frac{1}{2}\iint\left[\phi\ast(M+\tilde{M})\partial_{x}|M-\tilde{M}|+|\phi\ast(M-\tilde{M})||\partial_{x}M+\partial_{x}\tilde{M}|\right]g(x)h_{\delta}(t)\,dx\,dt\geq 0$
where $(x,t)$ is in place of $(\bar{x},\bar{t})$ to simplify the notation. Thanks to the choice of the functions $g(x)$ and $h_{\delta}(t)$ satisfying the properties above, and taking $\delta\to 0$, we get
@@
while we split the last integral of (3.46) into two terms to treat it. The first term is
@@
$\leq\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{s}^{t}\mathinner{\!\left\lVert M(\cdot,\tau)-\tilde{M}(\cdot,\tau)\right\rVert}_{L^{1}(\mathbb{R})}\,d\tau,$
@@
$\leq\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{\mathbb{R}}\partial_{y}M(y,\tau)+\partial_{y}\tilde{M}(y,\tau)\,dy=2\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}(\mathbb{R})}.$
@@
$\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}|\phi\ast(M-\tilde{M})|(\partial_{x}M+\partial_{x}\tilde{M})\,dx\,d\tau\leq\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}}\int_{s}^{t}\mathinner{\!\left\lVert M(\cdot,\tau)-\tilde{M}(\cdot,\tau)\right\rVert}_{L^{1}(\mathbb{R})}\,d\tau.$
@@
(3.47) $\mathinner{\!\left\lVert M(\cdot,t)-\tilde{M}(\cdot,t)\right\rVert}_{L^{1}(\mathbb{R})}$
$\leq\mathinner{\!\left\lVert M^{0}-\tilde{M}^{0}\right\rVert}_{L^{1}(\mathbb{R})}+t|A-\tilde{A}|_{\text{Lip}}+2\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{0}^{t}\mathinner{\!\left\lVert M(\cdot,\tau)-\tilde{M}(\cdot,\tau)\right\rVert}_{L^{1}(\mathbb{R})}\,d\tau.$
@@
$\leq\mathinner{\!\left\lvert\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}\omega\ast(\rho+\tilde{\rho})\partial_{x}|M-\tilde{M}|\,dx\,d\tau\right\rvert}\leq\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}|\omega\ast(\rho+\tilde{\rho})\mathinner{\!\left\lvert|\partial_{x}|M-\tilde{M}|\right\rvert}\,dx\,d\tau$
$\leq\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}2\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}\mathinner{\!\left\lvert\partial_{x}|M-\tilde{M}|\right\rvert}\,dx\,d\tau=\int_{s}^{t}\int_{\mathbb{R}}\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}|\partial_{x}(M-\tilde{M})|\,dx\,d\tau$
$\leq\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{s}^{t}\int_{\mathbb{R}}|\partial_{x}M|+|\partial_{x}\tilde{M}|\,dx\,d\tau=2\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}(t-s).$
@@
$\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}|\phi\ast(M-\tilde{M})|\partial_{x}(M+\tilde{M})\,dx\,d\tau$ $\leq\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}\int_{s}^{t}\int_{\mathbb{R}}\partial_{x}(M+\tilde{M})\,dx\,d\tau$
$=2\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}(t-s).$
@@
[…]
@@
[…]
$\hskip 70.0001pt\qquad||M^{0}-\tilde{M}^{0}||_{L^{1}(\mathbb{R})}+\left(|A-\tilde{A}|_{\text{Lip}}+4\mathinner{\!\left\lVert\omega\right\rVert}_{L^{\infty}(\mathbb{R})}\right)t\Big\}\,.$

### v2

@@ 3.2. The solution of the scalar balance law
for some given initial datum $M^{0}$. We have the following result.
@@
Assume (H) and (1.6). Consider the Cauchy problem (3.28) with $M^{0}$ a nondecreasing function, and suppose that there exists an $R^{0}>0$ such that
@@
    Given any $T>0$, the Cauchy problem (3.28) has a unique entropy weak solution
@@
    For any $T>0$, the entropy weak solution $M$ can be approximated by the discretized balance law (3.11) with initial datum (3.12), and thus by the sticky particle Cucker-Smale dynamics (2.1), i.e. there exists a sequence of initial data $M_{N}^{0}$ and fluxes $A_{N}$ satisfying the assumptions (3.14)-(3.13) respectively, such that the associated entropy weak solution $M_{N}$ constructed in Theorem 3.4 satisfies
@@
    Let $\tilde{M}$ be the entropy weak solution of
@@
    (3.31) $\|M(\cdot,t)-\tilde{M}(\cdot,t)\|_{L^{1}(\mathbb{R})}\leq e^{2t\|\phi\|_{L^{\infty}(\mathbb{R})}}\|M^{0}-\tilde{M}^{0}\|_{L^{1}(\mathbb{R})}+\frac{|A-\tilde{A}|_{Lip}}{2\|\phi\|_{L^{\infty}(\mathbb{R})}}(e^{t\|\phi\|_{L^{\infty}(\mathbb{R})}}-1),$
@@
    (3.32) $\|M(\cdot,t)-\tilde{M}(\cdot,t)\|_{L^{1}(\mathbb{R})}\leq\|M^{0}-\tilde{M}^{0}\|_{L^{1}(\mathbb{R})}+(|A-\tilde{A}|_{Lip}+4\|\omega\|_{L^{\infty}(\mathbb{R})})t.$
@@
Let $N\in\mathbb{N}$ be fixed, we construct two functions $M_{N}$ and $A_{N}$ satisfying the assumptions of Theorem 3.4. Let us consider positive masses $m_{i,N}$ such that
@@
Notice that $x_{i,N}^{0}\in[-R^{0},R^{0}],\,\forall i\in\{1,...,N\}$. Now, we can construct the functions $M_{N}^{0}$ and $A_{N}$ as in the following:
@@
These functions satisfy the hypotheses (3.14) and (3.13) of Theorem 3.4, respectively, then
@@
is an entropy weak solution to (3.11)-(3.12). We denote by $x_{i,N}(t)$, $i=1,\dots,N$ the trajectories of the particles involved in the Cucker-Smale dynamics with initial data $(m_{i,N},x_{i,N}^{0},v_{i,N}^{0})_{i=1}^{N}$. The initial velocities can be obtained using the relation
@@
Moreover, the quantities $M_{N}^{0}$ and $A_{N}$ approximate the quantities $M^{0}$ and $A,$ i.e., the following inequalities hold true
(3.35) $\|M_{N}^{0}-M^{0}\|_{L^{1}(\mathbb{R})}$ $\leq 2R^{0}\max_{1\leq i\leq N}m_{i,N},$
@@
and these imply that $M_{N}^{0}-M^{0}\to 0$ in $L^{1}(\mathbb{R})$ and $A_{N}\to A$ uniformly as $N\to\infty$ using (3.33).
@@
  • To prove [(3.35)], we denote $x_{0,N}^{0}\mathrel{\mathop{\ordinarycolon}}=-R^{0}$ and we write
    $\|M_{N}^{0}$ $-M^{0}\|_{L^{1}(\mathbb{R})}=\sum_{j=1}^{N}\int_{x_{j-1,N}}^{x_{j,N}}|M_{N}^{0}(x)-M^{0}(x)|\,dx$
@@
Let $T>0$ be a fixed time, since $M_{N}(\cdot,t)$ is uniformly bounded and nondecreasing for any $t\in[0,T],$ we can apply the Helly’s Theorem stating that there exists a $L^{1}_{loc}(\mathbb{R})$-convergent subsequence $M_{N_{k}}(\cdot,t)$. Applying a diagonal argument, it is possible to get another subsequence convergent in $L^{1}_{loc}(\mathbb{R})$ for all rational $t\in[0,T]$, we denote again this subsequence as $M_{N_{k}}$ to simplify the notation and we denote with $M(\cdot,t)$ the limit of this subsequence. The aim is to extend this convergence from $L^{1}_{loc}(\mathbb{R})$ to $L^{1}(\mathbb{R})$ and also to extend the result to irrational times. From (2.6), we know $\{x_{i,N}(t)\}_{i=1}^{N}\subset[-R(T),R(T)]$ with $R(T)=R^{0}+T\tilde{M}$, thus
@@
and so $M_{N_{k}}(t)-M(t)\to 0$ in $L^{1}(\mathbb{R})$ for all rational times $t\in[0,T]$.
@@
Moreover, by (3.37), it holds
$\|\partial_{t}M_{N_{k}}(x,t)\|_{\mathcal{M}}\leq\tilde{M},\qquad\forall t\in[0,T],$
@@
and so we can conclude that $M\in BV(\mathbb{R}\times[0,T])$. Now, we have to verify if the limit function $M$ satisfies the entropy condition (3.7). We check this using the Kružkov entropy pairs $(\eta,q)$
(3.39) $\eta(M)=|M-\alpha|,\quad q(M)=\text{sgn}(M-\alpha)(A(M)-A(\alpha)).$
We know that
@@
is an entropy weak solution of the discretized scalar balance law (3.11); this means that it satisfies the entropy inequality with the entropy/entropy-flux pair $(\eta,q_{N})$, where $\eta(m)=|m-\alpha|$, $q_{N}(m)=\text{sgn}(m-\alpha)(A_{N}(m)-A_{N}(\alpha))$ and $\alpha\in[0,1]$
@@
Now, we consider the subsequence $M_{N_{k}}$ and we pass to the limit, to simplify the notation we use $M_{N}$ in place of $M_{N_{k}}$. For a fixed $t\in[0,T]$ we obtain
(3.41) $\|\eta(M_{N})-\eta(M)\|_{L^{1}(\mathbb{R})}$ $=\int_{\mathbb{R}}|\eta(M_{N})-\eta(M)|\,dx\leq|\eta|_{Lip}\int_{\mathbb{R}}|M_{N}-M|\,dx$
$=|\eta|_{Lip}\|M_{N}-M\|_{L^{1}(\mathbb{R})}\to 0,$
@@
About the last term, we start by showing that $\phi\ast M_{N}$ converges to $\phi\ast M$ as $N\to+\infty$, indeed, since $\phi\in L^{\infty}(\mathbb{R})$, we have
$\|\phi\ast M_{N}-\phi\ast M\|_{L^{\infty}(\mathbb{R})}\leq\|\phi\|_{L^{\infty}(\mathbb{R})}\|M_{N}-M\|_{L^{1}(\mathbb{R})}\to 0.$
@@
$\|\partial_{x}\eta(M_{N})\|_{\mathcal{M}(\mathbb{R})}\leq|\eta|_{Lip}$
and this is also true for $\partial_{x}(\eta(M))$. Now, we are able to prove the convergence of the last term in (3.40)
@@
$\leq$ $\int_{0}^{T}\|\varphi(t)\|_{L^{\infty}(\mathbb{R})}\|\phi\ast M_{N}-\phi\ast M\|_{L^{\infty}(\mathbb{R})}\|\partial_{x}(\eta(M_{N})\|_{\mathcal{M}}\,dt$
$+\left|\int_{0}^{T}\int_{\mathbb{R}}(\phi\ast M)\varphi[\partial_{x}(\eta(M_{N}))-\partial_{x}(\eta(M))]\,dx\,dt\right|,$
and since $\|\phi\ast M_{N}-\phi\ast M\|_{L^{\infty}(\mathbb{R})}\to 0$, the first term on the right-hand side above goes to zero. On the second term we have
$\left|\int_{0}^{T}\int_{\mathbb{R}}(\phi\ast M)\varphi[\partial_{x}(\eta(M_{N}))-\partial_{x}(\eta(M))]\,dx\,dt\right|$
$\qquad\qquad=\left|\int_{0}^{T}\int_{\mathbb{R}}-\partial_{x}[(\phi\ast M)\varphi][(\eta(M_{N}))-(\eta(M))]\,dx\,dt\right|$
$\qquad\qquad\leq\int_{0}^{T}\int_{\mathbb{R}}\mathinner{\!\left\lvert-\partial_{x}[(\phi\ast M)\varphi][(\eta(M_{N}))-(\eta(M))]\right\rvert}\,dx\,dt$
$\qquad\qquad\leq\|\partial_{x}[(\phi\ast M)\varphi]\|_{L^{\infty}(\mathbb{R})}\int_{0}^{T}\|\eta(M_{N})-\eta(M)\|_{L^{1}(\mathbb{R})}\,dt\,.$
Since $\eta$ is Lipschitz continuous and by means of (3.38), the right-hand side above converges to zero and thus, $M$ is an entropy weak solution of (3.28).
@@
$+\iint\text{sgn}(M(x,t)-\tilde{M}(y,s))(A(M(x,t))-A(\tilde{M}(y,s)))\partial_{x}w(x,t,y,s)\,dx\,dt$
$+\iint(\phi\ast M)(x,t)(\partial_{x}|M(x,t)-\tilde{M}(y,s)|)\,dx\,dt.$
We perform the same for the entropy inequality of $\tilde{M}$, but with $\tilde{A}$ in place of $A$ and with $\alpha=\tilde{M}(y,s)$. We integrate both the inequalities above over the remaining free variables in both cases and adding them, we get
@@
$\iint\left[|M(x,t)-\tilde{M}(x,t)|g(x)h^{\prime}_{\delta}(t)+\text{sgn}(M-\tilde{M})(A_{+}(M)-A_{+}(\tilde{M}))g^{\prime}(x)h_{\delta}(t)\right]\,dxdt$
@@
$+\frac{1}{2}\iint\left[\phi\ast(M+\tilde{M})\partial_{x}|M-\tilde{M}|+|\phi\ast(M-\tilde{M})\|\partial_{x}M+\partial_{x}\tilde{M}|\right]g(x)h_{\delta}(t)\,dx\,dt\geq 0$
where $(x,t)$ is in place of $(\bar{x},\bar{t})$ to simplify the notation. Thanks to the previous choice of the functions $g(x)$ and $h_{\delta}(t)$, and taking $\delta\to 0$, we get
@@
while we split the last integral of (3.47) into two terms to treat it, the first one is
@@
$\leq\|\phi\|_{L^{\infty}(\mathbb{R})}\int_{s}^{t}\|M(\cdot,\tau)-\tilde{M}(\cdot,\tau)\|_{L^{1}(\mathbb{R})}\,d\tau,$
@@
$\leq\|\phi\|_{L^{\infty}(\mathbb{R})}\int_{\mathbb{R}}\partial_{y}M(y,\tau)+\partial_{y}\tilde{M}(y,\tau)\,dy=2\|\phi\|_{L^{\infty}(\mathbb{R})}.$
@@
$\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}|\phi\ast(M-\tilde{M})|(\partial_{x}M+\partial_{x}\tilde{M})\,dx\,d\tau\leq\|\phi\|_{L^{\infty}}\int_{s}^{t}\|M(\cdot,\tau)-\tilde{M}(\cdot,\tau)\|_{L^{1}(\mathbb{R})}\,d\tau.$
@@
(3.48) $\|M(\cdot,t)-\tilde{M}(\cdot,t)\|_{L^{1}(\mathbb{R})}$
$\leq\|M^{0}-\tilde{M}^{0}\|_{L^{1}(\mathbb{R})}+t|A-\tilde{A}|_{\text{Lip}}+2\|\phi\|_{L^{\infty}(\mathbb{R})}\int_{0}^{t}\|M(\cdot,\tau)-\tilde{M}(\cdot,\tau)\|_{L^{1}(\mathbb{R})}\,d\tau.$
@@
$\leq\mathinner{\!\left\lvert\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}\omega\ast(\rho+\tilde{\rho})\partial_{x}|M-\tilde{M}|\,dx\,d\tau\right\rvert}\leq\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}|\omega\ast(\rho+\tilde{\rho})|\mathinner{\!\left\lvert|\partial_{x}|M-\tilde{M}|\right\rvert}\,dx\,d\tau$
$\leq\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}2\|\omega\|_{L^{\infty}(\mathbb{R})}\mathinner{\!\left\lvert\partial_{x}|M-\tilde{M}|\right\rvert}\,dx\,d\tau=\int_{s}^{t}\int_{\mathbb{R}}\|\omega\|_{L^{\infty}(\mathbb{R})}|\partial_{x}(M-\tilde{M})|\,dx\,d\tau$
$\leq\|\omega\|_{L^{\infty}(\mathbb{R})}\int_{s}^{t}\int_{\mathbb{R}}|\partial_{x}M|+|\partial_{x}\tilde{M}|\,dx\,d\tau=2\|\omega\|_{L^{\infty}(\mathbb{R})}(t-s).$
@@
$\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}}|\phi\ast(M-\tilde{M})|\partial_{x}(M+\tilde{M})\,dx\,d\tau$ $\leq\|\omega\|_{L^{\infty}(\mathbb{R})}\int_{s}^{t}\int_{\mathbb{R}}\partial_{x}(M+\tilde{M})\,dx\,d\tau$
$=2\|\omega\|_{L^{\infty}(\mathbb{R})}(t-s).$
@@
(3.49) $\|M(\cdot,t)-\tilde{M}(\cdot,t)\|_{L^{1}(\mathbb{R})}\leq\|M^{0}-\tilde{M}^{0}\|_{L^{1}(\mathbb{R})}+(|A-\tilde{A}|_{\text{Lip}}+4\|\omega\|_{L^{\infty}(\mathbb{R})})t,$
@@
$\|M(\cdot,t)-\tilde{M}(\cdot,t)\|_{L^{1}(\mathbb{R})}$
$\leq\min\Big\{e^{2t\|\phi\|_{L^{\infty}(\mathbb{R})}}\|M^{0}-\tilde{M}^{0}\|_{L^{1}(\mathbb{R})}+\frac{|A-\tilde{A}|_{\text{Lip}}}{2\|\phi\|_{L^{\infty}(\mathbb{R})}}(e^{t\|\phi\|_{L^{\infty}(\mathbb{R})}}-1),$
$\hskip 70.0001pt\qquad\|M^{0}-\tilde{M}^{0}\|_{L^{1}(\mathbb{R})}+\left(|A-\tilde{A}|_{\text{Lip}}+4\|\omega\|_{L^{\infty}(\mathbb{R})}\right)t\Big\}\,.$